OpenStudy (anonymous):

There are 20 rabbits on an island. After six months the number of rabbits had increased to 100. If the number of rabbits increased exponentially, then how many rabbits will there be at the end of one year?

OpenStudy (dumbcow):

\[P = 20*b^t\] this exponential function represents the population after "t" months we know that \[20b^6 = 100\] or \[b^6 = 5\] to find b^12 we could square both sides (b^6)^2 = b^12 \[b^{12} = 25\] so population after 1 year is: \[P = 20b^{12} = 20*25 = 500\]

OpenStudy (anonymous):

thankyou so much